OpenStudy (anonymous):

Manuel builds and sells wooden crates. He made a table to show how much his crates cost. Which equation represents the situation? Manuel’s Crates Crates 2 4 6 8 Cost 25 45 65 85 A.y – 2 = −10(x – 25) B. y – 4 = 10(x – 45) C. y – 65 = −10(x – 6) D. y – 85 = 10(x – 8)

OpenStudy (anonymous):

@shaik0124 @EmmanuelRocksYou @

OpenStudy (anonymous):

Im sorry, @Briinndyy_69 i dont know how to do this :(

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@Emilywiggins6

OpenStudy (anonymous):

@roz_mar23 You know??

OpenStudy (anonymous):

A B C OR D

OpenStudy (anonymous):

Crates Cost 2 25 4 45 6 65 8 85 (X1 Y1) (X2 Y2) 4 45 8 85 Slope=m=rise/run---> Y2 -Y1= 85-45=40 X2-X1= 8-4=4 Y-Y1=,(X-X1) Pug in values Y-4=-10(X-45) B

OpenStudy (anonymous):

Sorry took so long...let me know if u get it right